# CATS: Enhancing Multivariate Time Series Forecasting by Constructing Auxiliary Time Series as Exogenous Variables

### v1

@@ 4 Experiments and Results
Comprehensive experiments were conducted on nine commonly-used MTSF datasets (details in Appendix A), including ETT (Zhou et al., 2021), Traffic, Electricity, Weather, Exchange Rate (Lai et al., 2018), and Multi (Lu et al., 2023). We compared CATS against baseline models from six recent MTSF studies: ARM (Lu et al., 2023), PatchTST (Nie et al., 2022), DLinear (Zeng et al., 2023), FedFormer (Zhou et al., 2022), Autoformer (Wu et al., 2021), and Informer (Zhou et al., 2021). We use the same experiment environment as these models for fair comparison. In CATS, a structurally simple two-layer MLP (2L) was employed as the time series predictor to demonstrate CATS’s capability of empowering univariate models to effectively learn inter-series relationships. We use a total of 8 ATS constructors, detailed in Table 6. We trained CATS with a combination of Mean Squared Error (MSE) and continuity loss, using the Adam optimizer (Kingma & Ba, 2015). Also, we applied common training strategies from recent studies, including last-value demeaning and random dropping (Kim et al., 2022; Zeng et al., 2023; Lu et al., 2023). To identify the enhancement specific from CATS, these strategies were also applied across all baseline models in subsequent ablation studies (§4.1). Models from Informer to PatchTST required tuning of the input length $L_{I}$, selecting from $L_{I}\in\{96,192,336,720\}$ for optimal results. However, CATS’s temporal sparsity module enables dynamic input cutoff for each channel, allowing a fixed $L_{I}=720$. In the main text, the performance metrics used included averaged MSE and calculated average MSE percentage compared to a pivotal model on the test set. For original MSE and Mean Absolute Error (MAE) data, please refer to the Appendix D. Details on hyper-parameter settings and implementations can be found in Appendix C. Some visualization examples of ATS and OTS results can be found in Figure 7.
@@
Table 1: Summary of MTSF results with forecasting horizons $L_{P}\in\{96,192,336,720\}$. See Table 7 for original results. Averages of MSE results (in parentheses) for each model on each dataset and MSE percentages (above parentheses) relative to DLinear are provided. The best and second best results are in bold and underlined, respectively. Average percentages (Avg%) across all datasets and average rankings (AvgRank) of each model, along with the count of first-place rankings (#Win), are also included.
Models CATS (2L) ARM PatchTST DLinear FedFormer Autoformer Informer
Electricity 89.6% 90.2% 95.6% 100.0% 128.4% 136.5% 186.9%
@@
Extended MTSF Experiment Results   In §1.1, we showed a shifting problem, which is introduced by ARM (Lu et al., 2023) to create the Multi dataset in our benchmarks to test models in strong multivariate relationship scenarios, where CATS demonstrated superior performance. To further assess CATS in more complex conditions with a greater number of series than 8 in Multi, we created the MultiX datasets. These datasets derive each series from a master AR(1) series, with each subsequent series lagging the previous by a set stride, adding and accumulating random noise in each shifting to challenge the model’s ability to learn generative dependencies rather than just relying on the first series. Eventually, the final series is noised offset from the first by a total of $S$ steps. We set $S=L_{I}=720$. If the number of extracted series, X, is 100, the dataset is named Multi100, and so forth. More details about this dataset are in Appendix A.
@@
Models CATS (2L) ARM PatchTST DLinear FedFormer Autoformer Informer
Multi20 67.0% 74.0% 109.7% 100.0% 230.7% 230.3% 338.3%
@@
Table 2 shows CATS’s exceptional performance in MultiX scenarios. CATS significantly outperforms previous models in average MSE and achieves the best results in 12 out of 16 sub-experiments. The original results in Table 8 reveals that CATS markedly surpasses previous models when $L_{P}\in\{96,192,336\}$, where true inter-series relationships are more prevalent. It’s noteworthy that when forecasting horizon $L_{P}$ extends to 720, CATS’s performance aligns more closely with traditional models. This pattern can be attributed to the reduced presence of actual inter-series dependencies in the input series for longer $L_{P}$ values. In essence, as the forecasting horizon expands, the influence of true inter-series relationships diminishes, leading to a scenario resembling a univariate AR(1) prediction with noise. In such cases, CATS’s channel sparsity feature becomes more prominent, effectively reducing the ATS’s influence and reverting the model to a simpler form of a univariate MLP. This behavior demonstrates CATS’s adaptability in handling varying strengths of multivariate relationships.
@@ 4.1 Ablation Studies
Time Series Predictors + CATS   In Table 3, we apply the CATS structure to a univariate model (DLinear), a multivariate model (Autoformer), and three basic univariate predictors (2L, IndLin, Mean). IndLin represents the use of independent linear layers for each ATS/OTS prediction. Mean fills the corresponding $L_{P}$ region for each ATS/OTS with its mean value of the input.
The integration of CATS significantly enhances the performance of each baseline. Examining the original results in Table 9, we find that CATS provides significant and stable performance enhancements in scenarios with strong multivariate relationships and effectively suppresses overfitting when these relationships are weaker, ensuring better results than the baselines. These enhancements are independent of the predictor’s structure. Even when the predictor is merely performing an average pooling (Mean), CATS is still able to transfer the inter-series relationships into the OTS output, surpassing the performance of most baselines. Note that IndLin performs the best in the sub-experiments with very strong inter-series dependencies, as shown in Table 9, which aligns with the example in §1.1. This highlights the powerful and flexible foundational fitting ability provided by the 2D attention-like structure for multivariate modeling.
@@ B.1 Continuity
[…]
where $f$ is the frequency. Let’s consider an ATS signal $A(t)$ generated based on $X(t)$. The continuity loss reduces the differences between adjacent points in ATS. Consequently, compared to $X(t)$, it imposes a penalty on the high-frequency components. This can be viewed as applying a low-pass filter $H(f)$ to ATS. Therefore, the autocorrelation function of ATS, $R_{A}(\tau)=R_{X}(\tau)*h(\tau)$, is influenced by this filtering, where *** denotes convolution and $h(\tau)$ is the inverse Fourier transform of $H(f)$. Here, the spectral density of $A(t)$ is
[…]
@@
For a random process driven by Brownian motion ($B_{t}$), the quadratic variation $[B]_{a}^{b}$ over the interval $[a,b]$ is equal to the length of the interval $b-a$. This indicates that the quadratic variation of Brownian motion is closely related to the volatility along its path. According to its fundamental theorem, the quadratic variation of any differentiable function over any interval is zero, reflecting the smooth nature of its trajectory. Consider an ATS $A(t)$, derived from an original time series $X(t)$. If $X(t)$ is driven by Brownian motion, then its quadratic variation over a certain interval is non-zero. However, by applying the Continuity principle, we aim to reduce the quadratic variation of $A(t)$. Let $A(t)$ be a continuous differentiable function, then its quadratic variation $[A]_{a}^{b}$ over the interval $[a,b]$ is zero, which implies:
$[A]_{a}^{b}=\lim_{\|P\|\to 0}\sum_{i=1}^{n-1}(A(t_{i+1})-A(t_{i}))^{2}=0$
@@ Appendix D Additional Experiment Results
Figure 7: Visualization of CATS (2L) results for the Multi and ETTh2 datasets with one example datapoint for each plot. 20 selected ATS ($A^{11}$ to $A^{30}$) and all the OTS are displayed. The red line represents the ground truth for each OTS channel. The blue line shows the predictor input for each ATS/OTS channel. The yellow line indicates the ATS/OTS prediction. The green line is the final forecasting results, combining ATS and OTS predictions. In the input domain, we use a green box to highlight the area remained after the temporal sparsity cutoff. Note that many input channels are turned into horizontal lines, which is caused by channel sparsity suppressing them; these series still have output prediction due to the bias in the 2L MLP. Full visualization can be found in supplementary materials.

### v2

@@ 4 Experiments and Results
Comprehensive experiments were conducted on nine commonly-used MTSF datasets (details in Appendix A), including ETT (Zhou et al., 2021), Traffic, Electricity, Weather, Exchange Rate (Lai et al., 2018), and Multi (Lu et al., 2024). We compared CATS against baseline models from six recent MTSF studies: ARM (Lu et al., 2024), PatchTST (Nie et al., 2022), DLinear (Zeng et al., 2023), FedFormer (Zhou et al., 2022), Autoformer (Wu et al., 2021), and Informer (Zhou et al., 2021). We use the same experiment environment as these models for fair comparison. In CATS, a structurally simple two-layer MLP (2L) was employed as the time series predictor to demonstrate CATS’s capability of empowering univariate models to effectively learn inter-series relationships. We use a total of 8 ATS constructors, detailed in Table 6. We trained CATS with a combination of Mean Squared Error (MSE) and continuity loss, using the Adam optimizer (Kingma and Ba, 2015). Also, we applied common training strategies from recent studies, including last-value demeaning and random dropping (Kim et al., 2022; Zeng et al., 2023; Lu et al., 2024). To identify the enhancement specific from CATS, these strategies were also applied across all baseline models in subsequent ablation studies (§4.1). Models from Informer to PatchTST required tuning of the input length $L_{I}$, selecting from $L_{I}\in\{96,192,336,720\}$ for optimal results. However, CATS’s temporal sparsity module enables dynamic input cutoff for each channel, allowing a fixed $L_{I}=720$. In the main text, the performance metrics used included averaged MSE and calculated average MSE percentage compared to a pivotal model on the test set. For original MSE and Mean Absolute Error (MAE) data, please refer to the Appendix D. Details on hyper-parameter settings and implementations can be found in Appendix C. Some visualization examples of ATS and OTS results can be found in Figure 7.
@@
Table 1: Summary of MTSF results with forecasting horizons $L_{P}\in\{96,192,336,720\}$. See Table 7 for original results. Averages of MSE results (in parentheses) for each model on each dataset and MSE percentages (above parentheses) relative to Repeat are provided. The best and second best results are in bold and underlined, respectively. Average percentages (Avg%) across all datasets and average rankings (AvgRank) of each model, along with the count of first-place rankings (#Win), are also included.
Models CATS (2L) ARM PatchTST DLinear Autoformer Informer Repeat
Electricity 9.24% 9.31% 9.87% 10.3% 14.1% 19.3% 100.0%
@@
Extended MTSF Experiment Results In §3.1, we showed a shifting problem, which is introduced by ARM (Lu et al., 2024) to create the Multi dataset in our benchmarks to test models in strong multivariate relationship scenarios, where CATS demonstrated superior performance. To further assess CATS in more complex conditions with a greater number of series than 8 in Multi, we created the MultiX datasets. These datasets derive each series from a master AR(1) series, with each subsequent series lagging the previous by a set stride, adding and accumulating random noise in each shifting to challenge the model’s ability to learn generative dependencies rather than just relying on the first series. Eventually, the final series is noised offset from the first by a total of $S$ steps. We set $S=L_{I}=720$. If the number of extracted series, X, is 100, the dataset is named Multi100, and so forth. More details about this dataset are in Appendix A.
@@
Models CATS (2L) ARM PatchTST DLinear Autoformer Informer Repeat
Multi20 58.2% 64.2% 95.3% 86.8% 200.0% 293.8% 100.0%
@@
Table 2 shows CATS’s exceptional performance in MultiX scenarios. CATS significantly outperforms previous models in average MSE and achieves the best results in 12 out of 16 sub-experiments. As the forecasting horizon expands, the influence of true inter-series relationships diminishes, leading to a scenario resembling a univariate AR(1) prediction with noise. In such cases, CATS’s channel sparsity feature becomes more prominent, effectively reducing the ATS’s influence and reverting the model to a simpler form of a univariate MLP. This behavior demonstrates CATS’s adaptability in handling varying strengths of multivariate relationships. We further discuss the importance of using MultiX as a benchmark dataset in §A.0.1.
@@ 4.1 Ablation Studies
Time Series Predictors + CATS In Table 3, we apply the CATS structure to a univariate model (DLinear), a multivariate model (Autoformer), and three basic univariate predictors (2L, IndLin, Mean). IndLin represents the use of independent linear layers for each ATS/OTS first-stage prediction. Mean fills the corresponding $L_{P}$ region for each ATS/OTS with its mean value of the input.
The integration of CATS significantly enhances the performance of each baseline. CATS provides significant and stable performance enhancements in scenarios with strong multivariate relationships and effectively suppresses overfitting when these relationships are weaker, ensuring better results than the baselines. These enhancements are independent of the predictor’s structure. Even when the predictor is merely performing an average pooling (Mean), CATS is still able to transfer the inter-series relationships into the OTS output, surpassing the performance of most baselines. Note that IndLin performs the best in the sub-experiments with very strong inter-series dependencies, as shown in Table 9, which aligns with the example in §3.1. This highlights the powerful and flexible foundational fitting ability provided by the 2D attention-like structure for multivariate modeling.
@@ 5 Conclusion, Future Works, and Limitation
Impact Statement
This paper advances Machine Learning by improving the accuracy and efficiency of multivariate time series forecasting with the proposed model, CATS. The method has beneficial applications like enhanced decision-making in domains such as finance and healthcare. While the societal impacts of this work are primarily positive, careful consideration and management of these technologies in sensitive applications are advised to prevent potential negative consequences.
@@ Appendix A Datasets
A.0.1 Why we need MultiX dataset?
Recent MTSF research has seen univariate models like PatchTST and DLinear achieve SOTA results in MTSF datasets, suggesting (a) potential design flaws in previous multivariate models and (b) possibly a lack of strong multivariate relationships in the current MTSF benchmark datasets. These univariate models inherently ignore inter-series/contextual information during inference, making them naturally more suited for datasets without inter-series relationships. However, their success is not sufficient to reject the existence or significance of contextual relationships in MTSF.
We believe CATS’s performance improvement over other models stems from its effective modeling of inter-series relationships. We used the Multi/MultiX datasets, which emphasize inter-series relationships, to demonstrate CATS’s effectiveness. However, since many datasets (like the ETTs) indeed lack significant inter-series dependencies, performance improvements from effectively modeling contextual information may appear marginal. We aim to highlight this through the CATS paper to prevent future MTSF research from misunderstanding due to benchmark dataset selection bias.
Among the commonly used benchmark datasets, ETTh1, ETTh2, ETTm1, ETTm2, Weather, and Exchange exhibit weak inter-series relationships. The Electricity dataset stands out as the only one that has more apparent inter-series dependencies. If a model enhances multivariate modeling capacities but only shows significant improvements on the Electricity dataset due to the lack of inter-series dependencies in others, it could be misleadingly deemed incremental upon evaluation. This seems unfair and misaligned with our actual expectations for MTSF research.
@@ B.1 Continuity
$S_{X}(f)=\int_{-\infty}^{\infty}R_{X}(\tau)e^{-i2\pi f\tau}d\tau,$
where $f$ is the frequency. Let’s consider an ATS signal $A(t)$ generated based on $X(t)$. The continuity loss reduces the differences between adjacent points in ATS. Consequently, compared to $X(t)$, it imposes a penalty on the high-frequency components. This can be viewed as applying a low-pass filter $H(f)$ to ATS. Therefore, the autocorrelation function of ATS, $R_{A}(\tau)=R_{X}(\tau)*h(\tau)$, is influenced by this filtering, where * denotes convolution and $h(\tau)$ is the inverse Fourier transform of $H(f)$. Here, the spectral density of $A(t)$ is
$S_{A}(f)=\int_{-\infty}^{\infty}R_{A}(\tau)e^{-i2\pi f\tau}d\tau=S_{X}(f)\cdot H(f).$
@@
For a random process driven by Brownian motion ($B_{t}$), the quadratic variation $B]_{a}^{b}$ over the interval $[a,b]$ is equal to the length of the interval bab-a. This indicates that the quadratic variation of Brownian motion is closely related to the volatility along its path. According to its fundamental theorem, the quadratic variation of any differentiable function over any interval is zero, reflecting the smooth nature of its trajectory. Consider an ATS $A(t)$, derived from an original time series $X(t)$. If $X(t)$ is driven by Brownian motion, then its quadratic variation over a certain interval is non-zero. However, by applying the Continuity principle, we aim to reduce the quadratic variation of $A(t)$. Let $A(t)$ be a continuous differentiable function, then its quadratic variation $A]_{a}^{b}$ over the interval $[a,b]$ is zero, which implies:
$A]_{a}^{b}=\lim_{\|P\|\to 0}\sum_{i=1}^{n-1}(A(t_{i+1})-A(t_{i}))^{2}=0$
@@ Appendix D Additional Experiment Results
Table 7 and Table 8 present full comprehensive MTSF results under various forecasting horizons, highlighting the performance of CATS under both general real-world and strongly inter-related series datasets. Table 9 shows the effectiveness of CATS with different time series predictors, while Table 10 focuses on the ablation studies of ATS principles and OTS shortcut. Additional performance comparisons of previous MTSF methods with and without CATS are detailed in Table 11. Table 12 and Table 13 provide insights into the contributions of different ATS constructors and the parameter proportions for each module within CATS, respectively, demonstrating the model’s robustness and efficiency.
Figure 7 provides the visualization of CATS (2L) results for the Multi and ETTh2 datasets. ATS from $A^{11}$ to $A^{30}$ and all the OTS are displayed. The red line represents the ground truth for each OTS channel. The blue line shows the predictor input for each ATS/OTS channel. The yellow line indicates the first-stage ATS/OTS prediction. The green line is the final forecasting results, combining ATS and OTS predictions. In the input domain, we use a green box to highlight the area remained after the temporal sparsity cutoff. Note that many input channels are turned into horizontal lines, which is caused by channel sparsity suppressing them; these series still have output prediction due to the bias in the 2L MLP.
@@
Table 11: Additional results of other previous MTSF methods with and without CATS. Input length $L_{I}$ are fixed as 720. The best and second best results are highlighted in bold and underlined, respectively.
Models 2L FITS FITS NBeats NBeats TiDE TiDE TimesNet TimesNet ARIMA Repeat
+CATS +CATS +CATS +CATS +CATS
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh2 (96) 0.259 0.327 0.271 0.336 0.269 0.340 0.643 0.560 0.300 0.363 0.273 0.339 0.270 0.338 0.388 0.422 0.315 0.381 0.392 0.407 0.432 0.422
ETTh2 (192) 0.309 0.368 0.330 0.374 0.322 0.381 1.226 0.738 0.387 0.420 0.334 0.378 0.329 0.388 0.407 0.430 0.394 0.432 0.503 0.467 0.534 0.473
ETTh2 (336) 0.329 0.383 0.353 0.395 0.349 0.401 1.347 0.787 0.424 0.448 0.356 0.400 0.351 0.405 0.399 0.438 0.391 0.435 0.572 0.512 0.591 0.508
ETTh2 (720) 0.358 0.415 0.378 0.422 0.375 0.430 1.411 0.839 0.436 0.469 0.385 0.431 0.380 0.431 0.425 0.456 0.350 0.407 0.622 0.541 0.588 0.517
Multi (96) 0.022 0.092 0.072 0.198 0.052 0.155 0.131 0.273 0.020 0.091 0.073 0.198 0.064 0.187 0.050 0.162 0.039 0.135 0.080 0.200 0.068 0.189
Multi (192) 0.055 0.146 0.153 0.288 0.141 0.242 0.304 0.425 0.067 0.152 0.145 0.285 0.114 0.250 0.093 0.219 0.087 0.211 0.190 0.298 0.143 0.273
Multi (336) 0.155 0.282 0.280 0.386 0.203 0.331 0.669 0.652 0.146 0.256 0.269 0.380 0.193 0.319 0.218 0.339 0.173 0.302 0.384 0.415 0.264 0.369
Multi (720) 0.419 0.464 0.684 0.587 0.513 0.531 1.368 1.001 0.480 0.493 0.632 0.559 0.547 0.551 0.627 0.569 0.542 0.539 1.064 0.653 0.617 0.551
Table 12: Additional ablation studies of ATS constructors. ”Full” represents the default CATS configuration (2L), with multiple baseline models added for comparison after the Pure Conv ablation. Different ATS constructors may be suited to different datasets, and some datasets might not require certain ATS constructors (as shown in the table below for ETTh2’s Conv, and ETTm2’s Id and Emb). However, thanks to the sparsity module, there’s no need for manual experimentation to select the most appropriate ATS constructors. The incomplete CATS models are still enough to beat previous SOTAs.
Models Full w/o Conv w/o NOConv w/o Iconv w/o Lin w/o Id w/o Emb Pure Lin Pure Conv ARM PatchTST DLinear Autoformer Informer
ETTh2 (196) 0.310 0.310 0.317 0.314 0.315 0.316 0.315 0.317 0.323 0.327 0.341 0.383 0.426 3.792
ETTm2 (196) 0.218 0.219 0.219 0.221 0.220 0.218 0.218 0.220 0.224 0.218 0.223 0.224 0.281 0.533
Table 13: The proportion of parameter sizes for each module within CATS for the ETTh2 setting ($C=7$, $L_{I}=720$, $L_{P}=96$) and the Electricity setting ($C=321$, $L_{I}=720$, $L_{P}=96$), calculated using PyTorch ”.numel()” method.
Module Predictor (2L) $F^{[\mathtt{NOConv}]}$ Temporal Sparsity Channel Sparsity $F^{[\mathtt{Conv}]}$ $F^{[\mathtt{Emb}]}$ $F^{[\mathtt{IConv}]}$ Output Projection $F^{[\mathtt{Lin}]}$
Params (ETTh2) 2353056 63648 50469 50434 21312 2880 700 518 96
Params% (ETTh2) 92.53% 2.50% 1.98% 1.98% 0.84% 0.11% 0.03% 0.02% 0.00%
Params (Electricity) 2353056 1872288 276863 2208556 629216 2880 700 2716 2608
Params% (Electricity) 32.02% 25.48% 3.77% 30.05% 8.56% 0.04% 0.01% 0.04% 0.04%
@@
Figure 7: Visualization of CATS (2L) results for the Multi and ETTh2 datasets with one example datapoint for each plot.